# SPRINT Enables Interpretable and Ultra-Fast Virtual Screening against Thousands of Proteomes

### v1

@@ Appendix E Structural visualizations
We visualize the attention patterns of the attention pooling layer on the protein-ligand bound structure of several PDB IDs. We compare the attention patterns of SPRINT-sm models with ProtBert and SaProt trained on the MERGED dataset. We see across these diverse proteins and ligands that on average, the SaProt model attends to residues closer to the ligand, while the ProtBert models often attend to residues far from the binding site.
Head 1
Head 2
Head 3
Head 4
@@
Figure S3: Analyzing the attention of the ProtBert model on PDB ID 2X4Z. Each row is the ProtBert model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
Head 1
Head 2
Head 3
Head 4
@@
Figure S4: Analyzing the attention of the ProtBert-1to1 model (trained with 1:1 positive to negative ratio) on PDB ID 2X4Z. Each row is the ProtBert model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
Head 1
Head 2
Head 3
Head 4
@@
Figure S5: Analyzing the attention of the SaProt model on PDB ID 2X4Z. Each row is the SaProt model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
Head 1
Head 2
Head 3
Head 4
@@
Figure S6: Analyzing the attention of the SaProt-1to1 model (trained with 1:1 positive to negative ratio) on PDB ID 2X4Z. Each row is the SaProt model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.

### v2

@@ Appendix F Structural visualizations
We visualize the attention patterns of the attention pooling layer on the protein-ligand bound structure of several PDB IDs. We compare the attention patterns of SPRINT models with ProtBert and SaProt trained on the MERGED dataset. We see across these diverse proteins and ligands that on average, the SaProt model attends to residues closer to the ligand, while the ProtBert models often attend to residues far from the binding site.
@@
Figure 9: Analyzing the attention of the ProtBert model on PDB ID 2X4Z. Each row is the ProtBert model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
@@
Figure 10: Analyzing the attention of the ProtBert-1to1 model (trained with 1:1 positive to negative ratio) on PDB ID 2X4Z. Each row is the ProtBert model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
@@
Figure 11: Analyzing the attention of the SaProt model on PDB ID 2X4Z. Each row is the SaProt model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.
@@
Figure 12: Analyzing the attention of the SaProt-1to1 model (trained with 1:1 positive to negative ratio) on PDB ID 2X4Z. Each row is the SaProt model trained with different seed. Each column is a different attention head. Gradient from white to red indicates the attention weight, where white is no attention and red is max attention for that head. The ligand is shown in blue.